Click here to view the latest rover updates
Latest Spirit Update
Spirit Remains Silent at Troy
Spirit Status for sol 2621-2627
Release Date: 5/26/11
More than 1,300 commands were radiated to Spirit as part of the recovery effort in an attempt to elicit a response from the rover. No communication has been received from Spirit since Sol 2210 (March 22, 2010). The project concluded the Spirit recovery efforts on May 25, 2011. The remaining, pre-sequenced ultra-high frequency (UHF) relay passes scheduled for Spirit on board the Odyssey orbiter will complete on June 8, 2011.
Total odometry is unchanged at 7,730.50 meters (4.80 miles).
Latest Opportunity Update
Rover Suffers a Series of Resets This Week
Opportunity Status for sol 3752-3758
Release Date: 8/22/14
Opportunity is moving south along the west rim of Endeavour Crater heading towards 'Marathon Valley,' a putative location for abundant clay minerals.
More recently, the incidence of Flash memory-induced resets has increased. The rover experienced resets on Sols 3754, 3757 and 3758 (Aug. 15, 18, and 19, 2014), which stops the onboard master sequence. Because of the project's vigilance and timely actions, the impact of the resets on rover science and exploration has been minimized. But the increase reset rate is compelling expedited corrective action to the Flash memory issue.
On Sol 3752 (Aug. 13, 2014), Opportunity bumped just a few feet (over a meter) to a surface target, called 'Mt. Edgecumbe.' On the next sol, the robotic arm was used to collect a Microscopic Imager mosaic of the target, and then placed the Alpha Particle X-ray Spectrometer (APXS) on the same for multi-sol integration. The reset on Sol 3754 (Aug. 15, 2014) cut short the integration to a single sol. On Sol 3757 (Aug. 18, 2014), another reset occurred, but real-time action from mission controllers reactivated the rover's sequence and Opportunity was able to complete the planned drive, achieving about 157 feet (48 meters). Another reset happened on Sol 3758 (Aug. 19, 2014), suspending the remote sensing observation on that sol. The plan ahead to resume rover activities, including driving, until the reset problem can be corrected.
As of Sol 3758 (Aug. 19, 2014), the solar array energy production was 692 watt-hours with an atmospheric opacity (Tau) of 0.888 and a solar array dust factor of 0.788.
Total odometry is 25.28 miles (40.69 kilometers).